Circular No. 8284

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

IRAS 05436–0007 A report was received from J. W. McNeil, Paducah, KY, of the appear-

ance of a new nebula in a dense region of the Lynds 1630 cloud in Orion, and apparently associated with IRAS 05436–0007, on his unfiltered CCD images taken with a 7.6-cm refractor on Jan. 23 UT. The object, which is located at α = 5^{h}46^{m}14^{s}, δ = −0^{o}05′.8 (equinox 2000.0), was then of total mag about 15-16 (with his CCD camera’s sensitivity peaking at 575 nm), but it is not present on seven Digitized Sky Survey images from 1951 to 1991. B. Reipurth, University of Hawaii (UH), confirms that a faint optical counterpart to IRAS 05436–0007 has gone into outburst and has produced a large reflection nebulosity, based on preliminary examination of red broadband CCD images obtained with K. Meech at the UH 2.2-m telescope on Jan. 31. Reipurth adds that this is a very rare event, apparently similar to that involving IRAS 05380–0728 (cf. Reipurth and Bally 1986, Nature 320, 336). The outburst may be an EX-Lup-type or FU-Ori-type eruption, driven by a sudden increase of accretion through a circumstellar disk, and thus in urgent need of observation (see Herbig 1977, Ap.J. 217, 693; Lehmann et al. 1995, A.Ap. 300, L9; Hartmann and Kenyon 1996, ARAA 34, 207). Reipurth also notes that HH 22 is in the line-of-sight of this new nebula but is not physically involved with the nebula (Eislöffel and Mundt 1997, A.J. 114, 280).

SUPERNOVAE 2003lu AND 2003lv Further to IAUC 8281, L. Strolger and A. Riess report another ap-

parent supernova (designated SN 2003lu) on 12 images taken during 2003 Sept. 13–14, located at α = 3^{h}32^{m}36^{s}.12, δ = −27^{o}55′01′′.4 (equinox 2000.0), which is 0′′.4 east and 0′′.3 south of the center of a nearby galaxy. ACS zband magnitudes: Sept. 1.89 UT, [27.5; 13.62, 26.1.

Further to IAUC 8125, Riess reports the discovery of another apparent supernova (mag z = 23.2) on 2003 Apr. 4.67, located at α = 12^{h}37^{m}28^{s}.97, δ = +62^{o}11′27′′.8. SN 2003lv was observed on nine additional dates, with the resulting light curve strongly supporting its being a supernova.

COMET C/2001 Q4 (NEAT) Visual total-magnitude estimates: Jan. 25.47 UT, 8.5 (D. Seargent,

Cowra, N.S.W., 6×35 binoculars); Feb. 3.97, 8.5 (W. Souza, Santos, Brazil, 11×80 binoc.); 9.97, 8.0 (A. Amorim, Florianopolis, Brazil, 7×50 binoc.).

2004 February 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green